Can u show me working for
X/3 + 2 = 9

Answer by rm29924(97)

x/3 + 2 = 9
we want to get x by itself
so we subtract 2 from both sides
x/3 + 2 = 9
- 2 -2
x/3=7
to get rid of the 3 on the bottom we multipy both sides by 3
3(x/3) = 7(3)
3 is the same as 3/1
3/1(x/3) 3x/3 and 3 cancels out
x = 21
plug in the answer to check your work
x/3 + 2 = 9
21/3 + 2 = 9
21 divied by 3 = 7
7+2=9
9=9